Item 13 13 of 20 Part B Review Constants Periodic Table The power cube transformer for a portable CD player has an output oThe “power cube” transformer for a portable CD player has an output of 9.3 V and 800 mA (both rms) when plugged into a 120 V outlet.
Answer #1

We know that in transformers

Vp/Vs = is/ip

ip = Current in primary coil = is*(Vs/Vp)

Given that

is = 800 mA

Vs = 9.3 V

Vp = 120 V

So,

ip = 800*9.3/120

ip = 62 mA (this is rms current in primary coil)

Now since we need peak current in primary coil, So

ip_peak = ip*sqrt 2

ip_peak = 62*sqrt 2

ip_peak = 87.68 mA = 88 mA (In two significant figures)

Please Upvote.